19 May 2010

All hands on deck

Setting up equipment in the new gas laboratory
Ernesto and Thomas setting up equipment in the new gas laboratory.

All people in camp are now working in their different positions/jobs.
In two warm laboratories, people are crowding to setup all their equipment, in the science trench all processors were today going through all aspects of core processing and the first cores were processed. The drillers work in shifts, and the two problem solvers, Steff and Jakob, fixed things.

Things in camp are beginning to settle into a routine and there is a constant flow of people in and out of the main dome. Lou and Sverrir have been helping Mille from the dog sled expedition to pack down her gear. It is planned that Mille and her dogs will be picked up from NEEM on Friday.

What we have done today:
1. Drilling and logging. We have past the clear ice from Interstadial 9 and gone through
  Stadial 10. Some cloudy bands show wavy patterns (micro folds).
2. Setting up equipment in warm laboratories.
3. Packing down gear from the dog sled expedition.
4. Processing school and ice core processing.
5. Logging temperatures in drill trench and science trench.
6. Setting up water vapour sampling site.
7. Setup extra table in science trench.

Ad.1: Drillers report:
Six runs today produced 18.50 meters. Most runs suffered high current (>11 A) after
about 2 meters of drilling. An attempt to slightly increase the cutting pitch by about 0.5 mm also resulted in drilling at too high of a current. In the final run damage was observed in the pump. Steff will repair and replace in the morning.
Also we suspected that the hollow shaft was beginning to clog, so a new procedure will be introduced to completely melt and dry the entire hollow shaft overnight at the conclusion of each evening shift.

Drilled 18.40 meters. Driller's depth 1795.44 meters.

Ad.4: Processed 3 bags, last bag 3197, 1758.35 meters and cut SC piece from bag 2331 to bag 2355 for CFA measurements.

Ad.5: Temperature in: Drill trench -20°C, science trench -23°C, in new cave -26°C.

Weather: blue sky or high thin clouds, -18°C to -11°C, 1-5 knots from SE.
Visibility: Unrestricted. With almost no wind and sunshine it was almost t-shirt weather.

FL, J.P. Steffensen


← Previous entry   Next entry →